# Kinematic and dynamical origins of mean-$p_{T}$ fluctuations in heavy-ion collisions

### v1

@@ Abstract
Event-by-event fluctuations of the mean transverse momentum (mean-$p_{T}$) provide a sensitive probe of collective dynamics beyond single-particle spectra and anisotropic flow. In this Letter, we present a systematic comparison of mean-$p_{T}$ fluctuation measurements with calculations based on a Bayesian-calibrated multistage hydrodynamic framework. The experimental definitions employed by the STAR and ALICE Collaborations are implemented explicitly and found to yield consistent results within controlled limits. We study the centrality and beam-energy dependence of the observable, its sensitivity to key soft-sector ingredients, and the impact of the kinematic $p_{T}$ acceptance. By introducing scaled-$p_{T}$ cuts, we demonstrate that a part of the apparent energy dependence arises from kinematic projection effects, while the remaining trends reflect genuine collective dynamics. Our results establish mean-$p_{T}$ fluctuations as a nontrivial and independent validation of calibrated hydrodynamic descriptions of the quark–gluon plasma.
@@ 1 Introduction
In this Letter, we present a systematic comparison of mean-$p_{T}$ fluctuation measurements with predictions from a Bayesian-calibrated multistage hydrodynamic model at top RHIC and LHC energies JETSCAPE:2020shq , JETSCAPE:2020mzn . We implement the STAR and ALICE definitions explicitly, examine the centrality and beam-energy dependence of the observable, and study its sensitivity to key soft-sector model ingredients. We further investigate the role of the kinematic $p_{T}$ acceptance by comparing results obtained with experimental and scaled $p_{T}$ cuts, allowing us to disentangle genuine dynamical trends from kinematic projection effects.
@@ 3.1 Baseline comparison and sensitivity to soft kinematics
While the STAR- and ALICE-style model evaluations are mutually consistent, the model systematically underestimates the measured fluctuation magnitude when the experimental lower $p_{T}$ cut of $0.2$ GeV is used. A similar underestimation is observed in the Bayesian inference study at $\sqrt{s_{NN}}=2.76$ TeV, where mean-$p_{T}$ fluctuation data were included in the calibration, yet the MAP parameter set nevertheless yields systematically smaller fluctuation magnitudes for the experimental acceptance JETSCAPE:2020mzn . This indicates that the trend observed at 200 GeV reflects a systematic feature of the calibrated framework rather than an isolated discrepancy. To investigate the origin of this difference, we varied the lower edge of the acceptance window within the model while keeping the upper cut fixed. As illustrated in Fig. 1, lowering $p_{T}^{\min}$ leads to a noticeable increase of $\sqrt{C_{m}}/\langle p_{T}\rangle$ and improves the agreement with the data.
@@ 4 Conclusions
When comparing results across beam energies, we demonstrated that a substantial part of the apparent energy dependence of $R_{p_{T}}$ obtained under fixed experimental $p_{T}$ cuts can arise from kinematic projection effects. Introducing a scaled-cut prescription that aligns the acceptance relative to the characteristic soft momentum scale qualitatively modifies the observed energy dependence, emphasizing the need for acceptance-aware comparisons in RHIC-to-LHC studies and beam energy scan programs.

### v2

@@ Abstract
Event-by-event fluctuations of the mean transverse momentum (mean-$p_{T}$) provide a sensitive probe of collective dynamics beyond single-particle spectra and anisotropic flow. We present a systematic study of mean-$p_{T}$ fluctuation observables using a Bayesian-calibrated multistage hydrodynamic framework, including quantitative comparisons to RHIC measurements and model-based investigations of beam-energy and kinematic-acceptance effects. The experimental definitions employed by the STAR and ALICE Collaborations are implemented explicitly and found to yield consistent results within controlled limits. We study the centrality and beam-energy dependence of the observable, its sensitivity to key soft-sector ingredients, and the impact of the kinematic $p_{T}$ acceptance. By introducing scaled-$p_{T}$ cuts, we demonstrate that a part of the apparent energy dependence arises from kinematic projection effects, while the remaining trends reflect genuine collective dynamics. Our results establish mean-$p_{T}$ fluctuations as a nontrivial and independent validation of calibrated hydrodynamic descriptions of the quark–gluon plasma.
@@ 1 Introduction
In this Letter, we present a systematic study of mean-$p_{T}$ fluctuations using a Bayesian-calibrated multistage hydrodynamic model JETSCAPE:2020shq , JETSCAPE:2020mzn , including quantitative comparisons to RHIC measurements and predictive investigations of beam-energy and kinematic-acceptance effects at LHC energies. We implement the STAR and ALICE definitions explicitly, examine the centrality and beam-energy dependence of the observable, and study its sensitivity to key soft-sector model ingredients. We further investigate the role of the kinematic $p_{T}$ acceptance by comparing results obtained with experimental and scaled $p_{T}$ cuts, allowing us to disentangle genuine dynamical trends from kinematic projection effects.
@@ 3.1 Baseline comparison and sensitivity to soft kinematics
While the STAR- and ALICE-style model evaluations are mutually consistent, the model systematically underestimates the measured fluctuation magnitude when the experimental lower $p_{T}$ cut of $0.2$ GeV is used. A similar underestimation is observed in the Bayesian inference study at $\sqrt{s_{NN}}=2.76$ TeV, where mean-$p_{T}$ fluctuation data were included in the calibration, yet the MAP parameter set nevertheless yields systematically smaller fluctuation magnitudes for the experimental acceptance JETSCAPE:2020mzn . This indicates that the trend observed at 200 GeV reflects a systematic feature of the calibrated framework rather than an isolated discrepancy. [^2] Mean-$p_{T}$ fluctuation measurements at LHC energies are available from ALICE and are reported primarily as a function of charged-particle multiplicity rather than in wide centrality classes ALICE:2014gvd , ALICE:2024apz . Performing a direct model–data comparison within the present framework would therefore require a dedicated reanalysis using multiplicity-based event selection and narrow multiplicity bins, which in turn would demand substantially higher event statistics to control statistical fluctuations. Such an analysis is beyond the scope of the present work. To investigate the origin of this difference, we varied the lower edge of the acceptance window within the model while keeping the upper cut fixed. As illustrated in Fig. 1, lowering $p_{T}^{\min}$ leads to a noticeable increase of $\sqrt{C_{m}}/\langle p_{T}\rangle$ and improves the agreement with the data.
@@ 3.4 Species dependence and effective kinematic projection
Within this context, the present species-resolved results for light hadrons provide a controlled baseline for collective dynamics at kinetic freeze-out. Heavy-flavor particles offer valuable complementary information on the degree of thermalization and coupling between heavy quarks and the medium, and fluctuation observables could, in principle, provide additional sensitivity beyond average flow measurements; indeed, within the JETSCAPE framework, heavy-flavor dynamics—particularly for open heavy mesons—have been studied using dedicated transport modules JETSCAPE:2022hcb , JETSCAPE:2023ikg , with significant progress in describing heavy-flavor $R_{AA}$. A systematic extension of the present fluctuation analysis to heavy-flavor hadrons would therefore require a dedicated treatment of non-equilibrium production and transport effects and is beyond the scope of this work.
@@ 4 Conclusions
When comparing model results across beam energies, we demonstrated that a substantial part of the apparent energy dependence of $R_{p_{T}}$ obtained under fixed experimental $p_{T}$ cuts can arise from kinematic projection effects. Introducing a scaled-cut prescription that aligns the acceptance relative to the characteristic soft momentum scale qualitatively modifies the observed energy dependence, emphasizing the need for acceptance-aware comparisons in RHIC-to-LHC studies and beam energy scan programs.